OpenStudy (anonymous):

giving medals!!! A painter is going to paint the trim of a window located 18 feet from the ground. If the base of the ladder is placed 9 feet from the side of the house, how long does the ladder need to be to reach the window?

OpenStudy (johnweldon1993):

''Looks like a right triangle and we want to use the pythagorean theorem to solve for the hypotenuse

OpenStudy (anonymous):

a2 + b2=c2?

OpenStudy (johnweldon1993):

Correct

OpenStudy (anonymous):

405??

OpenStudy (juanpablojr):

18*18+9*9=c^2 square root the answer and bam should be right...

OpenStudy (johnweldon1993):

Well remember you want the square root of that...

OpenStudy (johnweldon1993):

\[\large c = \sqrt{a^2 + b^2}\]

OpenStudy (anonymous):

20.12?

OpenStudy (johnweldon1993):

There we go!

OpenStudy (anonymous):

but whats that in sq root

OpenStudy (johnweldon1993):

What do you mean? That would be our result \[\large 20.12 ft\]

OpenStudy (anonymous):

thats not one of my choices though

OpenStudy (anonymous):

a.9 b. sq root 9 over 5

OpenStudy (anonymous):

c.27

OpenStudy (anonymous):

d. 9 sq root 3

OpenStudy (johnweldon1993):

None of those are correct \[\large c = \sqrt{18^2 + 9^2} = \sqrt{324 + 81} = \sqrt{405} = 9\sqrt{5} \]

OpenStudy (anonymous):

thank you!

OpenStudy (aum):

BTW, if you multiply 9 * sqrt(5) you will get your earlier answer of 20.12 in decimal.

OpenStudy (anonymous):

thanks @aum